science student is riding on a flatcar of a train traveing along a straight horizontal track at a conszant speed of 12.3 m/s. The student throws a ball along a path that she judges to Nplecth iretantbe in ine with the track. The students professor who is standing on the ground nearby, observes the bail to rise vértičally AaTries 0/20 Post Discussion Send Feedback
Answer #1

Given,

v = 12.3 m/s ; theta = 33 deg ;

The x and y component of launch velocity will be:

vx = 12.3 cos33 = 10.32 m/s

vy = 12.3 sin33 = 6.69 m/s

t = vy/g = 6.69/9.8 = 0.68 s

dy = 1/2 at^2

dy = 0.5 x 9.8 x 0.68^2 = 2.27 m

Hence, d = 2.27 m